Presented below is information related to BLANK Corporation for the current year.

Beginning inventory $ 611,300

Purchases 1,493,300

Total goods available for sale $2,104,600

Sales revenue 2,505,000

Compute the ending inventory, assuming that (a) gross profit is 48% of sales, (b) gross profit is 60% of cost, (c) gross profit is 38% of sales, and (d) gross profit is 25% of cost. (Round ratios for computational purposes to 1 decimal place, e.g. 78.7% and final answers to 0 decimal places, e.g. 28,987.)

Ending Inventory

(a) Gross profit is 48% of sales $

b) Gross profit is 60% of cost $

(c) Gross profit is 38% of sales $

(d) Gross profit is 25% of cost $
Answer #1

Calculate ending inventory

Ending inventory = Goods available for sale-Cost of goods sold

Ending inventory
(a) Gross profit is 48% of sales (2104600-2505000*52%) = 802000
b) Gross profit is 60% of cost (2104600-1565625) = 538975
(c) Gross profit is 38% of sales (2104600-2505000*62%) = 551500
(d) Gross profit is 25% of cost (2104600-2004000) = 100600